# Self-Improving Robust Preference Optimization

### v3

@@ Abstract
Both online and offline RLHF methods such as PPO and DPO have been extremely successful in aligning AI with human preferences. Despite their success, the existing methods suffer from a fundamental problem that their optimal solution is highly task-dependent (i.e., not robust to out-of-distribution (OOD) tasks). Here we address this challenge by proposing Self-Improving Robust Preference Optimization (SRPO), a practical and mathematically principled offline RLHF framework that is completely robust to the changes in the task. The key idea of SRPO is to cast the problem of learning from human preferences as a self-improvement process, which can be mathematically expressed in terms of a min-max objective that aims at joint optimization of self-improvement policy and the generative policy in an adversarial fashion. The solution for this optimization problem is independent of the training task and thus it is robust to its changes. We then show that this objective can be re-expressed in the form of a non-adversarial offline loss which can be optimized using standard supervised optimization techniques at scale without any need for reward model and online inference. We show the effectiveness of SRPO in terms of AI Win-Rate (WR) against human (GOLD) completions. In particular, when SRPO is evaluated on the OOD XSUM dataset, it outperforms the celebrated DPO by a clear margin of $\mathbf{15\%}$ after $5$ self-revisions, achieving WR of $\mathbf{90}\%$.
@@
1 Introduction
Reinforcement Learning from Human Feedback (RLHF) (Christiano et al., 2017) has rapidly become a standard method to align Large Language Models (LLMs). One of the main practical issues that all the prominent existing RLHF methods (offline or online) (Ouyang et al., 2022; Rafailov et al., 2023; Azar et al., 2023; Zhao et al., 2023b; Ahmadian et al., 2024) encounter is that their optimal solution heavily depends on the training task in terms of the distribution used to generate the preference data (behavior policy) (Munos et al., 2023; Azar et al., 2023). This makes the existing RLHF methods prone to out-of-distribution (OOD) tasks (Li et al., 2024; Kirk et al., 2024) where the evaluation distribution is significantly different from that of the behavior policy. Also, whenever the base/SFT models significantly differ from the behavior policy, the dependency of the RLHF solutions on the behavior policy makes the preference dataset and reward model less useful (Gao et al., 2022) as RLHF may undo the SFT/pretraining.
To address this challenge, we introduce an alternative approach for aligning LLMs from human preferences based on more principled and robust foundations. Our goal is to find a solution that is robust to the changes in the preference dataset, meaning that changes in the distribution from which the completions are sampled do not affect the final outcome of learning significantly. To achieve this goal, we exploit the concept of self-improving (Huang et al., 2022; Bai et al., 2022) language models. By self-improving LLM we refer to a model capable of enhancing its outputs recursively with each inference iteration. Our Self-Improving Robust Preference Optimization (SRPO) consists of two back-to-back optimization processes:
[…]
@@ 1 Introduction
(Step 2) Robust Preference Optimization of Generative Model: The next step is to exploit the self-improvement policy learned in the previous step to learn a generative LLM, $\pi$. The key idea here is that the best generative policy can be identified as a policy that generates completions requiring minimal improvement using the optimal self-improvement policy $\pi_{\dagger}$ derived in step $1$. This goal can be achieved by minimizing the objective of step $1$ with respect to the generative policy for in-context completions, $y$. Similar to step $1$, this problem, in its KL-regularized form, can also be solved analytically in terms of the optimal improvement policy $\pi_{\dagger}$ and the optimal generative policy $\pi$. More significantly, we show that the solution for steps $1$ and $2$ can be estimated jointly through a single supervised direct preference optimization scheme using only a dataset of annotated pair-wise completions. Thus, one can solve both for the self-improvement policy $\pi_{\dagger}$ and $\pi$ by minimizing the supervised learning objective of SRPO. Unlike existing RLHF methods, this solution is independent of the behavior policy and is therefore robust to its changes.
@@
As using the self-improvement model in SRPO is a significant departure from the existing paradigm for RLHF, we provide a high-level motivation for it in Sec. 2. We then formalize our objective for SRPO in Sec. 3, allowing for the joint optimization of both $\pi_{\dagger}$ and $\pi$ by optimizing an adversarial min-max objective. In Sec. 4 we present our main algorithmic/mathematical contribution: we prove that the preference probability $p$ can be expressed in terms of the log-likelihoods of the optimal self-improvement policy $\pi_{\dagger}^{*}$ and the log-likelihoods of the optimal robust generative policy $\pi^{*}$. This theoretical finding is the key result for SRPO: solving this system of equations through least-squares regression provides us with the practical supervised SRPO objective that solves for both policy and robust generative policy through a single supervised objective without any need for reward model or online inference. Our key theoretical finding is similar to the main result of DPO (Rafailov et al., 2023) in that both express preference probabilities in terms of the optimal policy. However, DPO result only holds when preference probabilities conform to Bradly-Terry model (Bradley & Terry, 1952), whereas our key result is general as it holds across all preference models. In Sec. 5 we further illustrate our argument on the robustness of SRPO by providing an in-depth analysis of the solution of SRPO and other direct preference optimization methods. We also showcase/analyze the robustness of SRPO on a simple synthetic example. Finally in Sec. 7 we conduct large-scale experiments on training LLMs with SRPO both on in-distribution and OOD summarization tasks and we compare the results with those of standard baselines.
@@ 2 Learning Self-Improvement Policy Through Preference Optimization
The goal of this section is to provide some motivation on why learning self-improvement models through preference optimization can be useful for learning a good policy. First, we start by considering a more fundamental question:
@@
To answer this question, we notice that human preferences provide information on the relation between more-preferred and less-preferred completions. This information can be used to improve the less-preferred completions towards the more preferred completions. In other words, we can learn a model of alignment mechanics, the rules on how to improve the completions to better match human preferences. This is arguably a more natural learning task, considering human preferences, than directly learning the highest preferred completion by humans, which is the goal of standard RLHF methods. Note that the highest preferred answer is very unlikely to be in our completion dataset, especially when the space of possibilities is the entirety of human language, and in particular when the completions are generated from some LLM, which is subpar to humans. Instead, it is more natural to learn that given a query $x$ and a completion $y$ what would be the improved completion upon $y$, i.e., learn the model that aims at improving the output of LLM through a self-improvement process. In this case, if our model has captured the underlying rules of human preference, then it can use that to improve the subpar completions towards the best completions.
The existing self-improvement-based pipelines mostly rely on the in-context learning ability of pretrained/SFT LLMs (Bai et al., 2022; Wei et al., 2023). In the following, we show how the self-improvement policy, $\pi_{\dagger}$, can be optimally trained alongside the generative policy, $\pi$, using pair-wise preferences.
@@ 3 SRPO Objective
We start by introducing some notations required for establishing our theoretical results.
@@
Notations. Let $x$ and $y$ denote a context and a completion drawn from the space of all possible contexts $\mathcal{X}$ and all possible completions $\mathcal{Y}$, respectively. The large language model (LLM) is represented by the probability distribution (policy) $\pi$ where $\pi(y|x)$ denotes the probability of the completion $y$ given the context $x$. In the remainder of this article, we consider three variants of this base LLM, the trainable model $\pi_{\text{train}}$ (for which we use the short-hand notation $\pi$), the reference model $\pi_{\text{ref}}$ and the behavior model $\mu$ from which the completions in pair-wise preference dataset is sampled.
@@
We also introduce the self-improvement $\pi(y^{\prime}|y,x)$ as a model that using a context $x$ and in-context completion (thought) $y$ aims at improving $y$ to a better completion $y^{\prime}$. Similar to base LLM, we can define a reference model $\pi_{\text{ref}}(y^{\prime}|y,x)$ also for the self-improvement model. Let $\mathcal{D}=\{x,y_{1},y_{2}\}$ be a dataset of contexts and completions where $y_{1}$ and $y_{2}$ are drawn independently from $\mu(\cdot|x)$. We then present every pair $y_{1},y_{2}$ to human annotators who express preferences for one of the completions, denoted as $y_{w}\succ y_{l}$ where $y_{w}$ and $y_{l}$ denote the preferred and dis-preferred actions amongst $\{y_{1},y_{2}\}$ respectively. We then write true human preference $p(y_{1}\succ y_{2}|x)$ the probability of $y_{1}$ being preferred to $y_{2}$ knowing the context $x$. The probability comes from the randomness of the choice of the human we ask for their preference. So $p(y_{1}\succ y_{2}|x)=\mathop{\mathbb{E}}_{h}[\mathbb{I}\{h\mbox{ prefers }y_{1}\mbox{ to }y_{2}\mbox{ given }x\}]$, where the expectation is over humans $h$.
@@
Consider a reference policy $\pi_{\text{ref}}$, and a real positive regularisation parameter $\beta\in\mathbb{R}^{*}_{+}$. Then, we define the Self-Improving Robust Preference Optimisation objective (SRPO) for every context $x$ as
@@
$J^{*}(x)=\min_{\pi}\max_{\pi_{\dagger}}\mathop{\mathbb{E}}_{\begin{subarray}{c}y_{1}\sim\pi(.|x)\\ y_{2}\sim\pi_{\dagger}(\cdot|y_{1},x)\end{subarray}}\bigg{[}p(y_{2}\succ y_{1}|x)-\beta D_{\text{KL}}(\pi_{\dagger}||\pi_{\text{ref}}|y_{1},x)+\beta D_{\text{KL}}(\pi||\pi_{\text{ref}}|x)\bigg{]},$ (1)
with the KL-regularization terms are defined as: $D_{\text{KL}}(\pi_{\dagger}||\pi_{\text{ref}}|y_{1},x)=\text{KL}(\pi_{\dagger}(\cdot|y_{1},x)||\pi_{\text{ref}}(\cdot|y_{1},x))$ and $D_{\text{KL}}(\pi||\pi_{\text{ref}}|x)=\text{KL}(\pi(\cdot|x)||\pi_{\text{ref}}(\cdot|x))$.
@@
In nutshell, this objective aims at (i) finding the best self-improvement policy $\pi^{*}_{\dagger}$ that takes every $y_{1}\sim\pi$ and improves it optimally with respect to the preference distribution $p$, i.e., the improved policy is most preferred to $y_{1}$, while keeping $\pi^{*}_{\dagger}$ close to the reference policy $\pi_{\text{ref}}$, (ii) minimizing the same objective to find the best (robust) policy $\pi^{*}$ for which the generated completions can be only minimally improved by the optimal self-improvement model $\pi^{*}_{\dagger}$. the min-max nature of this objective guarantees that self-improvement is effective for all policies (close to $\pi_{\text{ref}}$) as we are optimizing $\pi_{\dagger}$ in the worst-case scenario.
@@ 4 Offline Solution for Optimizing SRPO Objective
The optimization problem of Eq. (1) is a non-trivial optimization problem that often requires solving a two-stage adversarial optimization problem through the game-theoretic approaches, which are often challenging and difficult to scale up, (see e.g., Munos et al., 2023; Rosset et al., 2024; Calandriello et al., 2024, for how we can use game-theoretic approaches/objectives to train LLMs). Here, inspired by Rafailov et al. (2023); Azar et al. (2023), we aim at casting this complex optimization objective as a standard supervised learning problem that can be solved at scale given an offline pairwise preference dataset. To derive a practical algorithm for SRPO we first notice that the inner-maximization in the objective function of Eq. (1) can be solved in closed form as follows:
$\pi^{*}_{\dagger}(y_{2}|y_{1},x)=\frac{\exp\left(\frac{p(y_{2}\succ y_{1}|x)}{\beta}\right)\pi_{\text{ref}}(y_{2}|y_{1},x)}{Z^{*}(y_{1},x)},$ (2)
where $Z^{*}(y_{1},x)$ is the normalization factor. One can easily show that by plugging $\pi^{*}_{\dagger}$ in the objective function of Eq. (1) we obtain:
$J^{*}(x)=\min_{\pi}\mathop{\mathbb{E}}_{y_{1}\sim\pi(.|x)}\left[\beta(\log(Z^{*}(y_{1},x))+D_{\text{KL}}(\pi||\pi_{\text{ref}}|x))\right].$ (3)
Now by solving Eq. (2) with respect to $p(y_{2}\succ y_{1}|x)$ we obtain
$p(y_{2}\succ y_{1}|x)=\beta(\log(\pi^{*}_{\dagger}(y_{2}|y_{1},x))-\log(\pi_{\text{ref}}(y_{2}|y_{1},x))+\beta\log(Z^{*}(y_{1},x))).$ (4)
4.1 Optimizing the Self-Improvement Policy $\pi_{\dagger}$
We notice that using the convention $p(y_{1}\succ y_{1}|x)=\frac{1}{2}$ Eq. (4) implies
$\frac{1}{2}=\beta(\log(\pi^{*}_{\dagger}(y_{1}|y_{1},x))-\log(\pi_{\text{ref}}(y_{1}|y_{1},x)))+\beta\log(Z^{*}(y_{1},x)).$ (5)
Now by subtracting Eq. (4) from Eq. (5) we derive
$p(y_{2}\succ y_{1}|x)=\frac{1}{2}+\beta\left[\log\left(\frac{\pi^{*}_{\dagger}(y_{2}|y_{1},x)}{\pi_{\text{ref}}(y_{2}|y_{1},x))}\right)-\log\left(\frac{\pi^{*}_{\dagger}(y_{1}|y_{1},x)}{\pi_{\text{ref}}(y_{1}|y_{1},x)}\right)\right].$ (6)
This is our first key result that express preference $p(y_{2}\succ y_{1}|x)$ in terms of the optimal self-improvement policy $\pi^{*}_{\dagger}$. So we can enforce this equation for all $y_{1}$ and $y_{2}$ through following $\ell_{2}$ loss:
@@ 4.1 Optimizing the Self-Improvement Policy $\pi_{\dagger}$
$L_{\dagger}(\pi_{\dagger})=\mathop{\mathbb{E}}_{\begin{subarray}{c}y_{1},y_{2}\sim\mu(\cdot|x)\\ x\sim\rho\end{subarray}}\left[p(y_{2}\succ y_{1}|x)-\frac{1}{2}-\beta\left[\log\left(\frac{\pi_{\dagger}(y_{2}|y_{1},x)}{\pi_{\text{ref}}(y_{2}|y_{1},x))}\right)-\log\left(\frac{\pi_{\dagger}(y_{1}|y_{1},x)}{\pi_{\text{ref}}(y_{1}|y_{1},x)}\right)\right]\right]^{2}.$ (7)
Using the standard properties of $\ell_{2}$-norm to replace $P(y_{2}\succ y_{1}|x)$ with $\mathbf{1}(y_{1}\succ y_{2}|x)$, as $P(y_{2}\succ y_{1}|x)=\mathop{\mathbb{E}}(\mathbf{1}(y_{1}\succ y_{2}|x))$, in the objective of Eq. (15) allows us to derive the following sample loss for the improvement model:
$\widehat{L}_{\dagger}(\pi_{\dagger})=$ $\mathop{\mathbb{E}}_{\begin{subarray}{c}(y_{l},y_{w},x)\sim\mathcal{D}\end{subarray}}\left[\frac{1}{2}-\beta\left[\log\left(\frac{\pi_{\dagger}(y_{w}|y_{l},x)}{\pi_{\text{ref}}(y_{w}|y_{l},x))}\right)-\log\left(\frac{\pi_{\dagger}(y_{l}|y_{l},x)}{\pi_{\text{ref}}(y_{l}|y_{l},x)}\right)\right]\right]^{2}$ (8)
@@
$+\mathop{\mathbb{E}}_{\begin{subarray}{c}(y_{l},y_{w},x)\sim\mathcal{D}\end{subarray}}\left[\frac{1}{2}-\beta\left[\log\left(\frac{\pi_{\dagger}(y_{w}|y_{w},x)}{\pi_{\text{ref}}(y_{w}|y_{w},x))}\right)-\log\left(\frac{\pi_{\dagger}(y_{l}|y_{w},x)}{\pi_{\text{ref}}(y_{l}|y_{w},x)}\right)\right]\right]^{2}$
4.2 Optimizing the Robust Generative Policy $\pi$
In this section we want to derive an offline objective to optimize for generative model $\pi$ as well as improvement model $\pi_{\dagger}$. We start by collecting terms in Eq. (5) we obtain

### v4

@@ Abstract
Online and offline RLHF methods, such as PPO and DPO, have been highly successful in aligning AI with human preferences. Despite their success, however, these methods suffer from fundamental limitations: (a) Models trained with RLHF can learn from mistakes or negative examples through RL mechanism or contrastive loss during training. However, at inference time, they lack an innate self-improvement mechanism for error corrections. (b) The optimal solution of existing methods is highly task-dependent, making it difficult for them to generalize to new tasks. To address these challenges, we propose Self-Improving Robust Preference Optimization (SRPO), a practical and mathematically principled offline RLHF framework. The key idea behind SRPO is to cast the problem of learning from human preferences as a self-improvement process, mathematically formulated as a min-max objective that jointly optimizes a self-improvement policy and a generative policy in an adversarial fashion. Crucially, the solution for this optimization problem is independent of the training task, which makes it robust to its changes. We then show that this objective can be reformulated as a non-adversarial offline loss, which can be efficiently optimized using standard supervised learning techniques at scale. To demonstrate SRPO ’s effectiveness, we evaluate it using AI Win-Rate (WR) against human (GOLD) completions. When tested on the XSum dataset, SRPO outperforms DPO by a margin of $\mathbf{15\%}$ after $5$ self-revisions, achieving an impressive $\mathbf{90}\%$ WR. Moreover, on the challenging Arena-Hard prompts, SRPO outperforms both DPO and IPO (by $\mathbf{4\%}$ without revision and $\mathbf{6\%}$ after a single revision), reaching a $\mathbf{56\%}$ WR against against Llama-3.1-8B-Instruct.
1 Introduction
Reinforcement Learning from Human Feedback (RLHF) (Christiano et al., 2017) has quickly become a standard method to align large language models (LLMs). A key advantage of RLHF is that it allows the model to learn from its mistakes, allowing it to not only learn from good experiences but also from bad experiences and mistakes. Despite being able to learn from mistakes, the standard RLHF-tuned models are not equipped with an innate mechanism to self-improve and correct their mistakes at the time of inference (Kumar et al., 2024). If a model makes a critical mistake or omits crucial fact during inference, it struggles to recover. Another practical issue that all the prominent RLHF methods (offline or online) (Ouyang et al., 2022; Rafailov et al., 2023; Azar et al., 2023; Zhao et al., 2023b; Ahmadian et al., 2024) encounter is that their optimal solution heavily depends on the training task in terms of the distribution used to generate preference data (behavior policy) (Munos et al., 2023; Azar et al., 2023). This makes existing RLHF methods vulnerable to the cases in which the evaluation distribution of completions is significantly different from that of the behavior policy (Li et al., 2024b; Kirk et al., 2024).
Figure 1: Self-improvement by SRPO in terms of win rates against human (WR). We demonstrate robustness by training on TL;DR and evaluating on XSum. Gains on Direct Preference Optimization (DPO) are reported in text captions.
To address these challenges, we introduce an alternative approach for aligning LLMs from human preferences based on more principled and robust foundations. Our goal is to find a solution that is robust to the changes in the preference dataset, meaning that changes in the distribution from which the completions are sampled do not affect the final outcome of learning significantly. To achieve this goal, we exploit the concept of self-improving (Huang et al., 2022; Bai et al., 2022) language models. By self-improving LLM we refer to a model capable of enhancing its outputs recursively with each inference iteration. Our Self-Improving Robust Preference Optimization (SRPO) consists of two back-to-back optimization processes:
@@ 1 Introduction
(Step 2) Robust Preference Optimization of Generative Model: The next step is to exploit the self-improvement policy learned in the previous step to learn a generative LLM, $\pi$. The key idea here is that the best generative policy can be identified as a policy that generates completions requiring minimal improvement using the optimal self-improvement policy $\pi_{\dagger}$ derived in step $1$. This goal can be achieved by minimizing the objective of step $1$ with respect to the generative policy for in-context completions, $y$. Similarly to step $1$, this problem, in its KL-regularized form, can also be solved analytically in terms of the optimal improvement policy $\pi_{\dagger}$ and the optimal generative policy $\pi$. More significantly, we show that the solution for steps $1$ and $2$ can be estimated jointly through a single supervised direct preference optimization scheme using only a dataset of annotated pair-wise completions. Thus, one can solve both for the self-improvement policy $\pi_{\dagger}$ and $\pi$ by minimizing the supervised learning objective of SRPO. Unlike existing RLHF methods, this solution is independent of the behavior policy and is therefore robust to its changes.
@@
As using the self-improvement model in SRPO is a significant departure from the existing paradigm for RLHF, we provide a high-level motivation for it in Sec. 2. We then formalize our objective for SRPO in Sec. 3, allowing for the joint optimization of both $\pi_{\dagger}$ and $\pi$ by optimizing an adversarial min-max objective. In Sec. 4 we present our main algorithmic/mathematical contribution: we prove that the preference probability $p$ can be expressed in terms of the log-likelihoods of the optimal self-improvement policy $\pi_{\dagger}^{*}$ and the log-likelihoods of the optimal robust generative policy $\pi^{*}$. This theoretical finding is the key result for SRPO: solving this system of equations through least-squares regression provides us with the practical supervised SRPO objective that solves for both self-improvement policy and robust generative policy through a single supervised objective without any need for reward model or online inference. Our key theoretical finding is similar to the main result of DPO (Rafailov et al., 2023) in that both express preference probabilities in terms of the optimal generative policy. However, DPO result only holds when preference probabilities conform to Bradly-Terry model (Bradley & Terry, 1952), whereas our key result is general as it holds across all preference models. In Sec. 5 we further illustrate our argument on the robustness of SRPO by providing an in-depth analysis of the solution of SRPO and other direct preference optimization methods. We also showcase/analyze the robustness of SRPO on a simple synthetic example. Finally in Sec. 6 we conduct large-scale experiments on training LLMs with SRPO both on in-distribution and OOD summarization tasks and we compare the results with those of standard baselines.
@@ 2 Learning Self-Improvement Policy Through Preference Optimization
The goal of this section is to motivate why learning self-improvement models through preference optimization can be useful for training a GPT-style language model. One of the main challenges that affect the performance of sequence-to-sequence GPT-style models is that they are hard to recover from failure at the time of inference. This problem, which is known as exposure bias in the literature on sequence-to-sequence prediction (Xu et al., 2020; Arora et al., 2023; Bengio et al., 2015), is because at the time of inference, GPT-style models generate tokens step by step based on their own previous outputs, not the ground truth. This can lead to error accumulation: if the model generates an incorrect token early in the sequence (or a misleading token that may lead to an error later down the line), it may propagate errors since its subsequent predictions are based on the earlier incorrect (or misleading) token (Arora et al., 2023).
To address this shortcoming of GPT-style models, one may fine-tune the model such that it can revise its incorrect generations through a self-improvement/self-refinement process (Madaan et al., 2024; Hu et al., 2024; Huang et al., 2022; Bai et al., 2022). The simplest way to train such a self-improvement model is through supervised fine tuning (SFT) (Bai et al., 2022; Wei et al., 2023). To train a self-improvement model through SFT, one first need to build a supervised dataset of $(x,y,y^{*})$ in which $y$ is some initial completion than can be correct or incorrect and $y^{*}$ is the corrected version of $y$. The improvement model can then be fine-tuned using SFT to predict $y^{*}$ from the pair $(x,y)$.
However, training a SFT self-improvement model is a challenging task for the following reasons.
  1. 1.
    Creating a supervised dataset of $(x,y,y^{*})$ is resource intensive and hard to scale (Huang et al., 2022). Also, the existing standard SFT datasets are not created in this way.
  2. 2.
    Moreover, the SFT training pipeline for self-improvement can simply overfit to predict the correct answer $y^{*}$ only from the context $x$ and ignore the initial answer $y$ altogether or over-fit to specific refinement pattern (Li et al., 2024a). Thus it may not learn properly how to improve an incorrect $y$ to the correct $y^{*}$.
Alternatively, one might ask whether one can generalize the direct preference optimization methods (e.g., Rafailov et al., 2023; Azar et al., 2023) for learning self-improvement models using the standard human preference datasets. To answer this question we first try to answer a more fundamental question:
@@
To address this question, we observe that human preferences inherently encode valuable information about the relationship between more-preferred and less-preferred completions. This relational information can be leveraged to refine less-preferred completions, guiding them closer to those that better align with human preferences. In essence, this process involves learning a model of alignment mechanics—principles or "rules" that dictate how completions can be iteratively improved in terms of satisfying human preferences. This approach represents a more natural and intuitive learning task compared to directly predicting the highest-preferred completion, which is the focus of standard RLHF methods. Directly learning the optimal completion can be a challenging task, particularly when the completion space encompasses the full complexity of human language. Furthermore, the highest-preferred answer is unlikely to exist in the dataset, especially when completions are generated by large language models (LLMs) that inherently fall short of human-level quality.
Instead, it is more effective to model the improvement process itself: given a query $x$ and an initial completion $y$, the goal is to predict an improved completion $y^{\prime}$. This shift in perspective transforms the task into learning how to enhance subpar outputs from an LLM, rather than aiming to directly generate perfect responses. By capturing the underlying principles of human preference, the model can iteratively refine suboptimal completions $y$, progressively steering them towards the ideal outcome $y^{*}$. This framework not only aligns with the iterative nature of human reasoning but also facilitates continuous self-improvement, making it a robust and scalable approach for enhancing the performance of LLMs. In the following, we show how the self-improvement policy, $\pi_{\dagger}$, is trained alongside the generative policy, $\pi$, using pair-wise preferences.
@@ 3 SRPO Objective
We start by introducing some notations required to establish our theoretical results. Let $x$ and $y$ denote a context and a completion drawn from the space of all possible contexts $\mathcal{X}$ and all possible completions $\mathcal{Y}$, respectively. The large language model (LLM) is represented by the probability distribution (policy) $\pi$ where $\pi(y|x)$ denotes the probability of completion $y$ given the context $x$. In the remainder of this article, we consider three variants of this base LLM, the trainable model $\pi_{\text{train}}$ (for which we use the shorthand $\pi$), the reference model $\pi_{\text{ref}}$ and the behavior model $\mu$ from which the completions in pairwise preference dataset is sampled.
@@
We also introduce self-improvement $\pi(y^{\prime}|y,x)$ as a model that using a context $x$ and in-context completion (thought) $y$ aims at improving $y$ to better completion $y^{\prime}$. Similarly to the base LLM, we can define a reference model $\pi_{\text{ref}}(y^{\prime}|y,x)$ also for the self-improvement model. Let $\mathcal{D}=\{(x,y,y^{\prime})\}$ be a dataset of contexts and completions where $y$ and $y^{\prime}$ are drawn independently of $\mu(\cdot|x)$. We then present every pair $y,y^{\prime}$ to human annotators who express preferences for one of the completions, denoted as $y_{w}\succ y_{l}$ where $y_{w}$ and $y_{l}$ denote the preferred and dis-preferred actions amongst $\{y,y^{\prime}\}$ respectively. We then write the true human preference $p(y^{\prime}\succ y|x)$ the probability that $y^{\prime}$ is preferred to $y$ knowing the context $x$. The probability comes from the randomness of the choice of the human we ask for their preference. So $p(y^{\prime}\succ y|x)=\mathop{\mathbb{E}}_{h}[\mathbb{I}\{\{\}h\mbox{ prefers}y^{\prime}\mbox{ to }y\mbox{ given }x\}]$, where the expectation is over humans $h$.
@@
Consider a reference policy $\pi_{\text{ref}}$, and a real positive regularization parameter $\beta\in\mathbb{R}^{*}_{+}$. Then, we define the Self-Improving Robust Preference Optimization objective (SRPO) for every context $x$ as
@@
$J^{*}(x)=\min_{\pi}\max_{\pi_{\dagger}}\underset{\begin{subarray}{c}y\sim\pi(.|x)\\ y^{\prime}\sim\pi_{\dagger}(\cdot|y,x)\end{subarray}}{\mathop{\mathbb{E}}}\bigg{[}p(y^{\prime}\succ y|x)-\beta D_{\text{KL}}(\pi_{\dagger}||\pi_{\text{ref}}|y,x)+\beta D_{\text{KL}}(\pi||\pi_{\text{ref}}|x)\bigg{]},$ (1)
with the KL-regularization terms are defined as: $D_{\text{KL}}(\pi_{\dagger}||\pi_{\text{ref}}|y,x)=\text{KL}(\pi_{\dagger}(\cdot|y,x)||\pi_{\text{ref}}(\cdot|y,x))$ and $D_{\text{KL}}(\pi||\pi_{\text{ref}}|x)=\text{KL}(\pi(\cdot|x)||\pi_{\text{ref}}(\cdot|x))$.
@@
In nutshell, this objective aims at (i) finding the best self-improvement policy $\pi^{*}_{\dagger}$ that improves every $y\sim\pi$ optimally w.r.t. the preference distribution $p$, i.e., the improved policy is most preferred to $y$, while keeping $\pi^{*}_{\dagger}$ close to the reference policy $\pi_{\text{ref}}$, (ii) minimizing the same objective to find the best (robust) policy $\pi^{*}$ for which the generated completions can be only minimally improved by the optimal self-improvement model $\pi^{*}_{\dagger}$. The min-max nature of this objective makes self-improvement effective for all policies close to $\pi_{\text{ref}}$ as we optimize $\pi_{\dagger}$ in the worst-case scenario.
@@ 4 Offline Solution for Optimizing SRPO Objective
In this section, we show that the min-max objective of (1) can be transformed to a non-adversarial offline supervised loss. Thus it can be optimized at scale using standard optimization techniques.
4.1 Main Result
@@
The optimization problem of (1) is a non-trivial optimization problem that often requires solving a two-stage adversarial optimization problem through the game-theoretic approaches, which are often challenging and difficult to scale up, (see e.g., Munos et al., 2023; Rosset et al., 2024; Calandriello et al., 2024, for how we can use game-theoretic approaches/objectives to train LLMs). Here, inspired by Rafailov et al. (2023); Azar et al. (2023), we aim at casting this complex optimization objective as a standard supervised learning problem that can be solved at scale given an offline pairwise preference dataset. The main theoretical result of our work is then as follows
Theorem 1.
Given a context $x$ and the behavior policy $\mu(\cdot|x)$ let $\mu(y|x)$ and scalars $\alpha\in[0,1]$, $\beta>0$ we have that the solution of the min-max objective of (1) is obtained by minimizing the following loss
$L_{\alpha}(\pi,\pi_{\dagger})=$ $(1-\alpha)L(\pi,\pi_{\dagger})+\alpha L_{\dagger}(\pi_{\dagger}),$ (2)
where $L(\pi,\pi_{\dagger})$ and $L_{\dagger}(\pi_{\dagger})$ are defined respectively as follows
@@
$L_{\dagger}(\pi_{\dagger})=\mathop{\mathbb{E}}_{\begin{subarray}{c}y,y^{\prime}\sim\mu(\cdot|x)\\ x\sim\rho\end{subarray}}\left[p(y^{\prime}\succ y|x)-\frac{1}{2}-\beta\left[\log\left(\frac{\pi_{\dagger}(y^{\prime}|y,x)}{\pi_{\text{ref}}(y^{\prime}|y,x)}\right)-\log\left(\frac{\pi_{\dagger}(y|y,x)}{\pi_{\text{ref}}(y|y,x)}\right)\right]\right]^{2}.$ (3)
and
@@
$L(\pi,\pi_{\dagger})=$ $\mathop{\mathbb{E}}_{\begin{subarray}{c}y,y^{\prime}\sim\mu(\cdot|x)\\ x\sim\rho\end{subarray}}\Bigg{[}p(y^{\prime}\succ y|x)-\frac{1}{2}-\frac{\beta}{2}\bigg{[}\log\left(\frac{\pi_{\dagger}(y^{\prime}|y,x)}{\pi_{\text{ref}}(y^{\prime}|y,x)}\right)+\log\left(\frac{\pi(y|x)}{\pi_{\text{ref}}(y|x)}\right)$ (4)
$-\left(\log\left(\frac{\pi_{\dagger}(y|y^{\prime},x)}{\pi_{\text{ref}}(y|y^{\prime},x)}\right)+\log\left(\frac{\pi(y^{\prime}|x)}{\pi_{\text{ref}}(y^{\prime}|x)}\right)\right)\bigg{]}\Bigg{]}^{2}.$
4.2 Proof of Main Result
To prove the main result we first notice that the inner-maximization in the objective function of (1) is an instantiation of KL-regularized RL (Todorov, 2006). Thus it can be solved in analytical form and its solution is given by